# Bloch-Beilinson conjectures for Hecke characters and Eisenstein cohomology of Picard surfaces

### v2

@@ Abstract.
We consider certain families of Hecke characters $\phi$ over a quadratic imaginary field $F$. According to the Bloch-Beilinson conjectures, the order of vanishing of the $L$-function $L(\phi,s)$ at the central point $s=-1$ should be equal to the dimension of the space of extensions of the Tate motive ${\mathbb{Q}}(1)$ by the motive associated with $\phi$.
In this article, we construct candidates for the corresponding extensions of Hodge structures, assuming that the sign of the functional equation of $L(\phi,s)$ is $-1$. This is accomplished through the cohomology of variations of Hodge structures over Picard modular surfaces associated with $F$ and Harder’s theory of Eisenstein cohomology. Furthermore, we demonstrate that these extensions are naturally realized within certain biextensions. We outline a program to compute the biextension height and utilize it to establish the non-triviality of these extensions.
@@ 1. Introduction
Let $E$ be an elliptic curve over ${\mathbb{Q}}$ and denote by $r$ the rank of the finitely generated abelian group $E({\mathbb{Q}})$. Suppose that the sign $\epsilon(E)$ of the functional equation of the (completed) $L$-function of $E$ is $-1$. One of the striking consequences of the celebrated Gross-Zagier formula ([14]), in this case, is that if $L^{\prime}(E,1)\neq 0$, then $r\geq 1$. Since by the hypothesis on the sign, $L(E,s)$ vanishes at the central point $s=1$, this agrees with the Birch and Swinnerton-Dyer conjecture, which predicts in particular that $r$ should be precisely equal to the order of vanishing of $L(E,s)$ at $s=1$.
@@
One way to arrive at a reasonable prediction is by rephrasing the conjecture about elliptic curves in a form that lends itself to generalization. Given a rational point $P$ on an elliptic curve $E$ over ${\mathbb{Q}}$, one can consider the open subvariety $U:=E\setminus\{O,P\}$, where $O$ is the zero point. One looks at the localization exact sequence relating cohomology $H^{\bullet}_{Z}(E)$ of $E$ supported on $Z:=\{P,O\}$, cohomology with compact supports $H^{\bullet}_{c}(U)$ of $U$, and cohomology $H^{\bullet}(E)$. Since the divisor $P-O$ gives rise to a class in $H^{2}_{Z}(E)(1)$ that maps to 0 in $H^{2}(E)(1)$, one gets a morphism $\mathds{1}\rightarrow H^{2}_{Z}(E)(1)$ landing in the image of the connecting homomorphism from $H^{1}_{c}(U)(1)$. By pullback, this yields an exact sequence
$0\rightarrow H^{1}(E)(1)\rightarrow E_{P}\rightarrow\mathds{1}\rightarrow 0$
It turns out that the extension $E_{P}$ is non-trivial if and only if $P$ is of infinite order. This holds both in the sense of extensions of Galois representations (when taking étale cohomology) and of Hodge structures (when taking singular cohomology). Thus, one can reformulate the results of Gross-Zagier as saying that if $\epsilon(E)=-1$ and $L^{\prime}(E,1)\neq 0$, then there exists a non-trivial extension of $\mathds{1}$ by $H^{1}(E)(1)$ of geometric origin.
As explained in Section 3.3, the Bloch-Beilinson conjectures imply a generalization of the previous statement to the case of an algebraic Hecke character $\phi$ of $F$.
Conjecture 1.
Let $\phi$ be an algebraic Hecke character of $F$ of odd weight w. If $\epsilon(\phi)=-1$, then there exists a non-trivial extension
$E_{\phi}\in\mathop{\rm Ext}\nolimits^{1}_{\mathop{\rm MHS}\nolimits}(\mathds{1},H_{\phi}(\frac{w+1}{2}))$
@@
Here, $\mathop{\rm MHS}\nolimits$ is a suitable category of mixed Hodge structures with coefficients, to be defined in Section 2.9, and $H_{\phi}$ is the rank 1 pure Hodge structure associated with $\phi$. It arises as a subobject of the Hodge structure on the cohomology of a suitable abelian variety over $F$. The property of being of geometric origin will be made precise in Section 3.3; one can think of it as meaning “arising in the cohomology of an algebraic variety”.
The purpose of this paper is to initiate a program whose goal is confirming the above conjecture, for Hecke characters whose type and whose restriction $\phi_{{\mathbb{Q}}}$ to ${\mathbb{Q}}$ are of a special shape. Denote by $\omega_{F|{\mathbb{Q}}}$ the order two Dirichlet character associated to the extension $F|{\mathbb{Q}}$, and by $|\cdot|_{{\mathbb{I}}_{{\mathbb{Q}}}}$ the idelic norm. We prove the following result.
Theorem 1.
(Corollaries 7.2, 7.3) Let $k$ be a positive integer, and let $\phi$ be a Hecke character of $F$ of type $(k,-k-3)$ or $(-k-3,k)$, satisfying
$\phi_{{\mathbb{Q}}}=\omega_{F|{\mathbb{Q}}}|\cdot|^{3}_{{\mathbb{I}}_{{\mathbb{Q}}}}.$
@@
Note that our hypothesis on $\phi_{{\mathbb{Q}}}$ is equivalent to the very natural assumption that the conjugate inverse of $\phi$ verify
$\bar{\phi}^{-1}=\phi(-3)$
where $(-3)$ is the Tate twist (see Remark 6.1). [^1] We have chosen the formulation in the statement because it is the one more adapted to our approach using Eisenstein cohomology.
Work in progress of the two authors is aimed at showing that, at least when $L^{\prime}(\phi,-1)\neq 0$ (a “Gross-Zagier type” hypothesis), the extension provided by the above theorem is indeed non-trivial. In the rest of this introduction, we will first explain the principle of construction of our extension and the evidence for its non-triviality, then give an outline of the next steps of our program. Finally, we will survey the relation with other related works in the literature.
@@ 1.1. Ideas of construction and evidence for non-triviality
The strategy for proving our Theorem 1 is inspired by ideas of Harder ([18]) and makes use of his results in [17]. Starting from the Hecke character $\phi$, one can construct a subspace $\partial H^{2}(\phi)$ in degree 2 boundary cohomology, with coefficients in an appropriate local system, of a Picard modular surface $\mathrm{S}_{K}$, that is, a non-projective Shimura variety, attached to a ${\mathbb{Q}}$-algebraic group $\mathrm{G}$ such that $\mathrm{G}({\mathbb{R}})\simeq\mathop{\rm SU}\nolimits(2,1)$. Note that boundary cohomology of such a variety is computed in terms of the boundary of a compactification, but is independent of the choice of the latter. Then, one would like to lift elements of $\partial H^{2}(\phi)$ to classes in the cohomology of $\mathrm{S}_{K}$ itself. Through Langlands’ theory of Eisenstein operators, one sees that the desired lifts can be provided by classes represented by Eisenstein series, well defined as soon as their constant term converges. For Hecke characters whose infinity type and whose shape are as in our Theorem, this convergence is ensured precisely when $L(\phi,-1)=0$. Hence, in this case, we find that $\partial H^{2}(\phi)$ is a subspace of Eisenstein cohomology, i.e. of the restriction map $r$ from cohomology of $\mathrm{S}_{K}$ to its boundary cohomology.
A natural way to enforce the vanishing $L(\phi,-1)=0$ is to ask that $\epsilon(\phi)=-1$. Now, Rogawski’s study of the automorphic spectrum of unitary groups in three variables ([29, 30, 31]) shows that exactly under this sign hypothesis, a suitable subspace $H^{2}(\phi)_{!}$ can be constructed inside the interior cohomology of Picard surfaces with coefficients in the local system as before, where interior cohomology is defined as the kernel of the restriction map $r$. The subspace $H^{2}(\phi)_{!}$ is related to a $\mathrm{G}({\mathbb{A}}_{f})$-representation induced from $\phi$, and is non-trivial when $\phi$ has the shape prescribed by our Theorem.
@@
which is in fact an extension of Hodge structures. Classical results allow us to describe the Hodge structure on the piece on the left, while the one on the piece on the right can be computed thanks to the analysis of the degeneration of variations of Hodge structure to the boundary of the Baily-Borel compactification $\mathrm{S}^{*}_{K}$ of $\mathrm{S}_{K}$ of [1]. It turns out that up to a suitable twist, the above extension is nothing but an extension
$0\rightarrow H_{\phi}(-1)\rightarrow E_{0}\rightarrow\mathds{1}^{\oplus\dim\partial H^{2}(\phi)}\rightarrow 0$
@@
Note also that the two pieces in the above extension have been constructed from modules under the Hecke algebra which have repeating eigenvalues. The fact that the hypothesis $\epsilon(\phi)=-1$ implies the existence of a sub-Hecke module of interior cohomology of Picard surfaces, isomorphic at all finite places either to a induced representation from an Hecke character, or to a subquotient of it, is indeed one of the striking consequences of Rogawski’s work that we crucially exploit. We see this feature of our construction as the main reason to hope that, at least for a careful choice of the class in $\partial H^{2}(\phi)$, such an extension may be non split. Note that by contrast, the mixed Hodge structure on the first cohomology of a modular curve is split precisely because interior and Eisenstein cohomology have different eigenvalues under the action of the Hecke algebra.
@@ 1.2. Our program
This program explains the presence of subsection 7.2 of the present paper, devoted to the construction of the biextensions $B_{\phi}(\Psi,\Psi^{\prime})$. In order to make the computation of their biextension height accessible, one needs to provide explicit representatives of cohomology classes in $\partial H^{2}(\phi)$, by means of differential forms. The construction of such explicit representatives, using the language of Lie algebra cohomology, has recently been achieved in the work [2] by the two authors. The latter should hence be seen as providing one of the crucial ingredients required to carry out the above program. Note that it is precisely in the computation of the biextension height that the term $L^{\prime}(\phi,-1)$ is expected to appear, making its non-vanishing necessary (and sufficient) for the non-vanishing of the pairing itself, and hence for the non-vanishing of our extension class.
@@ 1.3. Related work and further future directions
When $\phi$ is an algebraic Hecke character of odd weight $w$, the Bloch-Beilinson conjectures ([8], [5]) are much more precise than their consequence studied in this paper, i.e. than Conjecture 1. Namely, they say that the order of vanishing of $L(\phi,s)$ at the central point $s=\frac{w+1}{2}$ should be equal to the (conjecturally finite) rank of a group of algebraic cycles, and they predict the value of the first non-trivial Taylor coefficient $L^{*}(\phi,\frac{w+1}{2})$ of $L(\phi,s)$ at the central point, up to an algebraic number.
Note that $s=\frac{w+1}{2}$ is a critical point, in the sense of Deligne [12], and hence there exists a canonical period associated to it as in op. cit.. Moreover, Deligne predicted that when non-zero, the value of $L(\phi,s)$ at such a point should be an algebraic multiple of his period. Deligne’s conjecture was later proven for algebraic Hecke characters of $CM$-fields by Blasius [6] and the recent preprint [21] proposes a proof for general number fields.
When the order of vanishing of $L(\phi,s)$ at the central point is greater than zero, the algebraic cycles whose existence is predicted by Bloch and Beilinson should give rise to non-trivial extensions, both as Hodge structures and Galois representations, through appropriate Abel-Jacobi maps. Moreover, the “transcendental part” of the value $L^{*}(\phi,\frac{w+1}{2})$ should be given by the product of Deligne’s period and of the determinant of an height pairing, and related to the extension class of the associated extension of Hodge structures in a precise way. In this paper, we propose candidates for the latter extensions. The program that we have outlined in the previous section of this introduction should moreover make it possible to compute their (conjecturally non-zero) extension class. Our further goal is of course to show that this class, and the height pairing that we plan to employ for its calculation, have indeed the expected relationship with the value $L^{*}(\phi,\frac{w+1}{2})$.
Galois representation-theoretic extensions have been constructed in [3], under the hypothesis of odd order of vanishing of $L(\phi,s)$ at the central point, for essentially the same kind of Hecke characters $\phi$ of $F$ that we consider here (the case of even order of vanishing is studied in [19]). As here, one crucial input in loc. cit. is represented by Rogawski’s results, and the source of the Galois representations considered in these works is again the cohomology of Picard modular surfaces. On the other hand, in order to obtain the desired extensions, the authors use $p$-adic families of automorphic forms, without looking at the geometry of the boundary of the Baily-Borel compactification and at Eisenstein cohomology. Hence, their extensions are not of geometric origin in our sense. It would be interesting to study the Galois-theoretic counterpart of our constructions, via étale cohomology, and to put it in relation to their methods.

### v4

@@ Abstract.
We consider certain families of Hecke characters $\phi$ over a quadratic imaginary field $F$. According to the Bloch-Beilinson conjectures, the order of vanishing of the $L$-function $L(\phi,s)$ at the central point s=1s=-1 should equal the dimension of the space of extensions of the Tate motive ${\mathbb{Q}}(1)$ by the motive associated with $\phi$.
In this article, we construct candidates for the corresponding extensions of Hodge structures, assuming that the sign of the functional equation of $L(\phi,s)$ is 1-1. This is accomplished through the cohomology of variations of Hodge structures over Picard modular surfaces associated with $F$ and Harder’s theory of Eisenstein cohomology.
Furthermore, we demonstrate that these extensions are naturally realized within certain biextensions and outline a program to establish their non-triviality.
@@ 1. Introduction
Let $E$ be an elliptic curve over $\mathbb{Q}$} and denote by $r$ the rank of the finitely generated abelian group $E({\mathbb{Q}})$. Suppose that the sign $\epsilon(E)$ of the functional equation of the (completed) $L$-function of $E$ is 1-1. One of the striking consequences of the celebrated Gross-Zagier formula ([20]), in this case, is that if $L^{\prime}(E,1)\neq 0$, then $r\geq 1$. Since by the hypothesis on the sign, $L(E,s)$ vanishes at the central point $s=1$, this agrees with the Birch and Swinnerton-Dyer conjecture, which predicts more precisely that $r$ should be equal to the order of vanishing of $L(E,s)$ at $s=1$.
@@
One way to arrive at a reasonable prediction is by rephrasing the conjecture about elliptic curves in a form that lends itself to generalization. Given a rational point $P$ on an elliptic curve $E$ over $\mathbb{Q}$}, one can consider the open subvariety $U:=E\setminus\{O,P\}$, where $O$ is the zero point. One looks at the localization exact sequence relating cohomology $H^{\bullet}_{Z}(E)$ of $E$ supported on $Z:=\{P,O\}$, cohomology with compact supports $H^{\bullet}_{c}(U)$ of $U$, and cohomology $H^{\bullet}(E)$. Since the divisor POP-O gives rise to a class in $H^{2}_{Z}(E)(1)$ that maps to 0 in $H^{2}(E)(1)$, one gets a morphism $\mathds{1}\rightarrow H^{2}_{Z}(E)(1)$ landing in the image of the connecting homomorphism from $H^{1}_{c}(U)(1)$. By pullback, this yields an exact sequence
$0\rightarrow H^{1}(E)(1)\rightarrow E_{P}\rightarrow\mathds{1}\rightarrow 0\,.$
It turns out that the extension $E_{P}$ is non-trivial if and only if $P$ is of infinite order (for more details, see the discussion in Remark 3.4). This holds both in the sense of extensions of Galois representations (when taking étale cohomology) and of Hodge structures (when taking singular cohomology). Thus, one can reformulate the results of Gross-Zagier as saying that if $\epsilon(E)=-1$ and $L^{\prime}(E,1)\neq 0$, then there exists a non-trivial extension of $\mathds{1}$ by $H^{1}(E)(1)$ of geometric origin, i.e. defined through a long exact sequence arising from the cohomology of an algebraic variety.
As explained in Subsection 3.4, the Bloch-Beilinson conjectures imply a generalization of the previous statement to the case of an algebraic Hecke character $\phi$ of $F$. To state it, let $\alpha$ denote both the non-trivial element of the Galois group of $F$ and the morphism that it induces on the idèles ${\mathbb{I}}_{F}$ of $F$. Denote by $|\cdot|_{{\mathbb{I}}_{F}}$ the idelic norm on ${\mathbb{I}}_{F}$, and for a Hecke character $\phi$ of $F$, define the Hecke character $\phi^{\perp}$ by putting
$\phi^{\perp}(x):=(\phi(\alpha(x)))^{-1},\ \ \ x\in{\mathbb{I}}_{F}\,.$
Conjecture 1 (Conjecture 3.5).
Let $\phi$ be an algebraic Hecke character of $F$ of odd weight w, such that $\phi^{\perp}=\phi|\cdot|_{{\mathbb{I}}_{F}}^{w}$. If the sign $\epsilon(\phi)$ of the functional equation of $\phi$ is 1-1, then there exists a non-trivial extension
$E_{\phi}\in\mathop{\rm Ext}\nolimits^{1}_{\mathop{\rm MHS}\nolimits}\left(\mathds{1},H_{\phi}\left(\frac{w+1}{2}\right)\right)$
@@
Here, $\mathop{\rm MHS}\nolimits$ is a category of mixed Hodge structures with coefficients, to be defined in Subsection 2.9, and $H_{\phi}$ is the rank 1 pure Hodge structure associated with $\phi$. It is a sub-object of the Hodge structure on the cohomology of a suitable abelian variety over $F$. The property of being of geometric origin will be made precise in Subsection 3.2; one can think of it as meaning “arising from the cohomology of an algebraic variety”. Note that under the hypothesis on $\phi^{\perp}$, there is indeed a well-defined sign $\epsilon(\phi)$, and the assumption $\epsilon(\phi)=-1$ implies that the $L$-function of $\phi$ vanishes at the central point $s=\frac{w+1}{2}$ (see Remark 3.2).
The purpose of this paper is to initiate a program whose goal is to confirm the above conjecture for certain Hecke characters of weight 3-3. We prove the following result.
Theorem 1 (Theorem 7.1.(3)).
Let $k$ be a positive integer, and let $\phi$ be a Hecke character of $F$ of type (k,k3)(k,-k-3) or (k3,k)(-k-3,k), satisfying [^1] Note the hypothesis on $\phi^{\perp}$ is equivalent to asking that the restriction $\phi_{{\mathbb{Q}}}$ of $\phi$ to $\mathbb{I}$$_{{\mathbb{Q}}}$ verifies $\phi_{{\mathbb{Q}}}=\omega_{F|{\mathbb{Q}}}|\cdot|^{3}_{{\mathbb{I}}_{{\mathbb{Q}}}}$ where $\omega_{F|{\mathbb{Q}}}$ is the order two Dirichlet character associated to the extension $F|{\mathbb{Q}}$ (see Remark 6.1 for this equivalence). This is the formulation that appears in the statement of Theorem 7.1. It is the one more adapted to our approach using Eisenstein cohomology.
(1) $\phi^{\perp}=\phi|\cdot|_{{\mathbb{I}}_{F}}^{-3}\,.$
@@
Work in progress of the two authors is aimed at showing that, at least when $L^{\prime}(\phi,-1)\neq 0$ (a “Gross-Zagier type” hypothesis), the extension provided by the above theorem is indeed nontrivial. In the rest of this introduction, we will first explain the principle of construction of our extension and the evidence for its non-triviality, then give an outline of the next steps of our program. Finally, we will survey the relation with other related works in the literature.
@@ 1.1. Ideas of construction and evidence for non-triviality
The strategy for proving Theorem 1 is inspired by ideas of Harder ([24]) and makes use of his results in [23]. Starting from the Hecke character $\phi$, one can construct a subspace $\partial H^{2}(\phi)$ in degree 2 boundary cohomology, with coefficients in an appropriate local system $V_{k}$, of a Picard modular surface $\mathrm{S}_{K}$, that is, a non-projective Shimura variety, attached to a $\mathbb{Q}$}-algebraic group $\mathrm{G}$ such that $\mathrm{G}({\mathbb{R}})\simeq\mathop{\rm SU}\nolimits(2,1)$. Note that the boundary cohomology of such a variety is computed in terms of the boundary of a compactification, but is independent of the choice of the latter. Then, one would like to lift elements of $\partial H^{2}(\phi)$ to classes in the cohomology of $\mathrm{S}_{K}$ itself. Through Langlands’ theory of Eisenstein operators, one sees that the desired lifts can be provided by classes represented by Eisenstein series, well defined as soon as their constant term converges. For Hecke characters satisfying the hypotheses of our Theorem, this convergence is ensured precisely when $L(\phi,-1)=0$. Hence, in this case, we find that $\partial H^{2}(\phi)$ is a subspace of Eisenstein cohomology, i.e. of the image of the restriction map
$r:H^{\bullet}(\mathrm{S}_{K},V_{k})\rightarrow\partial H^{\bullet}(\mathrm{S}_{K},V_{k})$
from $V_{k}$-valued cohomology of $\mathrm{S}_{K}$ to its $V_{k}$-valued boundary cohomology.
If $\phi$ is of weight 3-3 and satisfies (1), then a natural way to enforce the vanishing $L(\phi,-1)=0$ is to ask that $\epsilon(\phi)=-1$. Now, Rogawski’s study of the automorphic spectrum of unitary groups in three variables ([36, 37, 38]) shows that precisely when $\phi$ satisfies this sign hypothesis and (1), a suitable subspace $H^{2}(\phi)_{!}$ can be constructed inside the interior cohomology of Picard surfaces with coefficients in the local system $V_{k}$. Here, interior cohomology is defined as the kernel of the restriction map $r$, and the space $H^{2}(\phi)_{!}$ is related to a cuspidal $\mathrm{G}({\mathbb{A}}_{f})$-representation induced from $\phi$.
@@
which is in fact an extension of Hodge structures. Classical results allow us to describe the Hodge structure on the piece on the left, while the one on the piece on the right can be computed based on the analysis of the degeneration of variations of the Hodge structure to the boundary of Baily-Borel compactification $\mathrm{S}^{*}_{K}$ of $\mathrm{S}_{K}$ carried out in [2]. It turns out that up to a suitable twist, the above extension is nothing but an extension
$0\rightarrow H_{\phi}(-1)\rightarrow E_{0}\rightarrow\mathds{1}^{\oplus\dim\partial H^{2}(\phi)}\rightarrow 0\,.$
@@
The above extension is constructed starting from modules under the Hecke algebra with repeating eigenvalues. This happens because of a striking consequence of Rogawski’s work: the hypothesis $\epsilon(\phi)=-1$ implies the existence of a sub-Hecke module of interior cohomology of Picard surfaces, isomorphic at all finite places either to an induced representation from an Hecke character, or to a quotient of it. We see this feature of our construction as the main reason to hope that, at least for a careful choice of the class in $\partial H^{2}(\phi)$, such an extension may be non split. Note that by contrast, the mixed Hodge structure on the first cohomology of a modular curve is split, precisely because, in that case, interior and Eisenstein cohomology have different eigenvalues under the action of the Hecke algebra.
@@ 1.2. Our program
This program explains the presence of Section 8 of the present paper, devoted to the construction of the biextensions $B_{\phi}(\Psi,\Psi^{\prime})$. To make the computation of their biextension height accessible, one needs to provide explicit representatives of cohomology classes in $\partial H^{2}(\phi)$, by means of differential forms. The construction of such explicit representatives, using the language of Lie algebra cohomology, has recently been achieved in the work [5] by the two authors. The latter should hence be seen as providing one of the crucial ingredients required to complete the above program. Note that it is precisely in the computation of the biextension height that the term $L^{\prime}(\phi,-1)$ is expected to appear. Its non-vanishing should then be necessary (and sufficient) for the non-vanishing of the pairing itself, and hence for the non-vanishing of our extension class. This will be the subject of the forthcoming article [4].
@@ 1.3. Related work and further future directions
When $\phi$ is an algebraic Hecke character of odd weight $w$, the Bloch-Beilinson conjectures ([11], [8]) are much more precise than their consequence studied in this paper, i.e., than Conjecture 1. Namely, they say that the order of vanishing of $L(\phi,s)$ at the central point $s=\frac{w+1}{2}$ should be equal to the (conjecturally finite) rank of a group of algebraic cycles, and they predict the value of the first non-zero Taylor coefficient $L^{*}(\phi,\frac{w+1}{2})$ of $L(\phi,s)$ at the central point, up to an algebraic number.
Note that $s=\frac{w+1}{2}$ is a critical point, in the sense of Deligne [16], and hence there exists a canonical period associated to it as in op. cit.. Moreover, Deligne predicted that when non-zero, the value of $L(\phi,s)$ at such a point should be an algebraic multiple of his period. Deligne’s conjecture was later proven for algebraic Hecke characters of $CM$-fields by Blasius [9], and the recent preprint [27] proposes a proof for general number fields.
When the order of vanishing of $L(\phi,s)$ at the central point is greater than zero, the algebraic cycles whose existence is predicted by Bloch and Beilinson should give rise to non-trivial extensions, both as Hodge structures and Galois representations, through suitable Abel-Jacobi maps. Moreover, the “transcendental part” of the value $L^{*}(\phi,\frac{w+1}{2})$ should be given by the product of Deligne’s period and of the determinant of a height pairing. In this paper, we propose candidates for the expected extensions; moreover, the program that we have outlined in the previous section of this introduction should make it possible to compute their (conjecturally non-zero) extension class through a biextension height. Our further goal is, of course, to show that this biextension height is related to the aforementioned height pairing and that it bears the expected relationship with the value $L^{*}(\phi,\frac{w+1}{2})$.
Galois representation-theoretic extensions have been constructed in [6], under the hypothesis of odd order of vanishing of $L(\phi,s)$ at the central point, for essentially the same kind of Hecke characters $\phi$ of $F$ that we consider here (the case of even order of vanishing is studied in [25]). As here, one crucial input in loc. cit. is represented by Rogawski’s results, and the source of the Galois representations considered in these works is again the cohomology of Picard modular surfaces. On the other hand, to obtain the desired extensions, the authors use $p$-adic families of automorphic forms, without looking at the geometry of the boundary of the Baily-Borel compactification and at Eisenstein cohomology. Hence, their extensions are not of geometric origin in our sense. It would be interesting to study the Galois-theoretic counterpart of our constructions, via étale cohomology, and to put it in relation to their methods.